Derive the dimension of power


1
Expert's answer
2021-09-30T10:20:07-0400

By definition, the power is work per second. Thus, the units of power are:


"[P] = \dfrac{J}{s}"

Since "J = kg\cdot m^2/s^2", have:


"[P] = \dfrac{kg\cdot m^2}{s^3}"

The dimension is then:


"[P] = ML^2T^{-3}"

Answer. "[P] = ML^2T^{-3}".